OpenStudy (anonymous):

xy=32 2xy^2=32 find the value of p and q when x=2^p and y=2^q

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[xy=32\\ 2xy^2=32\] right?

OpenStudy (anonymous):

s

OpenStudy (anonymous):

ok so we can solve for \(x\) and \(y\) by dividing

OpenStudy (anonymous):

\[\frac{xy}{2xy^2}=\frac{32}{32}=1\] or \[\frac{1}{2y}=1\]

OpenStudy (anonymous):

that makes \(y=\frac{1}{2}\) substitute this back in to either equation to find \(x\)

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

x = 64

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[64=2^6,\frac{1}{2}=2^{-1}\]